Let X be a random variable from a distribution with density  g(x) = 3x^{2}cdot 1 [0,1] (x). Calculate the variance of random variable 3.22 +1
Answer #1

The probability density function of X is

g(z) = 32.2 , 0-2, < 1

Now,

5 5 0 0 0

and

0.428571

Hence,

Var(X) = E( X 4)-[E(X2)|2 = 0.428571-0.62 = 0.0685714

and Var(3X2 + 1) = Var(3X2) = 321 ar(X2) = 9 * 0.0685714 = 0.617143